Circular No. 7588

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2001 (87) 1 M. E. Brown and J. L. Margot, California Institute of Technology,

report the discovery of a satellite to (87) Sylvia. The projected separation between primary and secondary was 0′′.59 (1200 km) in p.a. 97^{o} on Hband images obtained on Feb. 18.5 UT using the adaptive-optics system at the 10-m W. M. Keck II telescope on Mauna Kea. Little relative motion was apparent in images obtained an hour apart (upper limit 0′′.02). The brightness ratio was measured to be 420 ± 70, implying a ∼ 1:20 ratio of diameters. Confirming observations were obtained at the Keck II telescope by I. de Pater and H. Roe on Feb. 19.4, 20.4, and 22.4, and the data indicate an orbital period on the order of 4 days.

JUPITER I (IO) I. de Pater, F. Marchis, and H. Roe, University of California at Berke-

ley; B. Macintosh, Lawrence Livermore National Laboratories; and S. Acton and D. Le Mignant, Keck Observatory, report an extreme volcanic outburst on Jupiter I: “We observed Io on Feb. 19, 20, and 22 UT with the Adaptive Optics system on the Keck II telescope in the bands J, H, and K′. On Feb. 20 and 22, we observed the hemisphere that approximately faced Jupiter. On Feb. 22, the region to the northeast of Loki was very bright in all three bands (K′ ∼ 5), being detectable even against the sunlit disk of Io. The hot spot is located at west longitude 334^{o} ± 3^{o} and north latitude 40^{o} ± 3^{o}, close to the position of Surt. This region did not show any discernible activity on Feb. 20. From the 3-band photometry, the temperature is estimated to be ∼ 1000–1100 K; an H-band spectrum taken at the same time suggests a temperature of ∼ 1700–1800 K. We encourage follow-up observations to determine the time evolution of this outburst.”

XTE J1543–568 R. Kaptein, BeppoSAX Science Operation Center, Rome, and Space

Research Organization Netherlands (SRON), Utrecht; J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and SRON; and J. Heise, SRON, write: “The Wide Field Cameras on BeppoSAX have for the first time detected XTE J1543–568. In an observation with unit 1 during Feb. 5.3-8.7 UT, the average flux was ∼ 9 mCrab (2–9 keV) at position α = 15^{h}44^{m}01^{s}, δ = −56^{o}45′.9 (equinox 2000.0; 99-percent-confidence error radius 2′.0). The position provides an improvement by a factor of 4 in accuracy (cf. IAUC 7369).”

2001 February 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green